OpenStudy (acal21):

[integrate from 0 to infinity] (ln(x)/x^2

ganeshie8 (ganeshie8):

did you try lnx = u ?

OpenStudy (acal21):

yes, it works , only problem is i am stuck in the part plugging in infinity to the equation \[(1/x)\ln(x)-1/x ] 0 and infinity

ganeshie8 (ganeshie8):

\[\large \int\limits_0^{\infty }\dfrac{\ln x }{x^2}~dx = \lim\limits_{a\to 0, b\to \infty} \int\limits_a^b\dfrac{\ln x }{x^2}~dx\]

ganeshie8 (ganeshie8):

now its a proper integral, evaluate it and take the limits

ganeshie8 (ganeshie8):

\[\large = \lim\limits_{a\to 0, b\to \infty} \left(\dfrac{\ln a+1}{a} - \dfrac{\ln b +1}{b}\right)\]

ganeshie8 (ganeshie8):

right ?

OpenStudy (acal21):

yes thanks